Two airplanes leave Toronto at the same time, one travelling west and the other east. Their average speeds differ by 5km/h. After 1 3/4 hours they were 980 km apart. Find the speed of each plane.
Answer by nerdybill(5411)

Two airplanes leave Toronto at the same time, one travelling west and the other east. Their average speeds differ by 5km/h. After 1 3/4 hours they were 980 km apart. Find the speed of each plane.
.
Let x = speed of one plane
then
x+5 = speed of other plane
.
1 3/4 = 1.75 hours
.
1.75x + 1.75(x+5) = 980
1.75x + 1.75x + 8.75 = 980
1.75x + 1.75x = 971.25
35x = 971.25
x = 277.5 km/hr
.
solution:
277.5 and 282.5 km/hr